OpenStudy (anonymous):

A pancake recipe requires one and one half cups of milk to one cup of flour. If three and one half cups of milk is used, what quantity of flour will be needed, according to the recipe?

OpenStudy (mathstudent55):

This is a proportion problem. You have a ratio between milk and flour given with the recipe. You want the ratio to be the same when you use a larger amount of milk. The ratio of milk to flour in the original recipe is milk/flour = 1.5/1 Now you're gouig to use 3.5 cups of milk to f cups of flour: 3.5/f now set the ratios equal to each other. That's a proportion. Then solve the proportion 1.5/1 = 3.5/f